In triangle ABC, the degree measure of angle A is 20 degree greater than the degree measure of angle B. Angle C is a right angle. What is the measure of angle B?
Since a triangle is 180 & c is 90 the other two sides get split therefore 45 but A is 20 degree more so B should be 25. Is this right

Answer by ankor@dixie-net.com(22740)

In triangle ABC, the degree measure of angle A is 20 degree greater than the degree measure of angle B. Angle C is a right angle. What is the measure of angle B?
As you said A + B = 90
A = (B+20)
therefore
(B+20) + B = 90
2B = 90 - 20
2B = 70
B = 70/2
B = 35 degrees
then
A = 35 + 20 = 55 degrees